OpenStudy (destinymasha):

A radio station had 250 tickets to a concert. They gave away 4 times as many tickets to listeners as to employees. How many tickets did they give away to employees? A. 50 B. 4 C. 5 D. 54

OpenStudy (mathstudent55):

Let e = number of tickets given to employees. Using the variable e, how many tickets did they give to listeners?

OpenStudy (destinymasha):

4 x as much as to employees?

OpenStudy (mathstudent55):

Right, so if the number of tickets given to employees is e, the number of tickets given to listeners is 4e. So far you follow?

OpenStudy (destinymasha):

yea

OpenStudy (mathstudent55):

Ok, now how many tickets did they give out altogether to the employees plus to the listeners, in terms of e and 4e?

OpenStudy (destinymasha):

250?

OpenStudy (mathstudent55):

That's the total they gave out. But my question is they gave out e tickets to employees and 4e tickets to listeners; the total they gave out in terms of e is e plus 4e. What is e plus 4e written with math symbols?

OpenStudy (mathstudent55):

plus in math is +

OpenStudy (destinymasha):

I think 4e ^2 or 4e + e

OpenStudy (destinymasha):

Sorry if i am learning a bit slow. [:

OpenStudy (mathstudent55):

The second one. 4e + e is the total number of tickets given out. But as you wrote above, the total number of tickets given out is 250, so 4e + e = 250 This is the equation you need to solve

OpenStudy (destinymasha):

ok,

OpenStudy (mathstudent55):

To solve 4e + e = 250, add the 4e and e on the left side. Remember that e is the same as 1e, so 4e + e = 4e + 1e = 5e So 4e + e = 250 becomes 5e = 250 Now divide both sides by 5 to get e

OpenStudy (destinymasha):

does e = 50?

OpenStudy (mathstudent55):

Correct. e = 50. Since we let e = the number of tickets given to employees, then the answer is There were 50 tickets given to employees.

OpenStudy (destinymasha):

yayyyyyy!!! Thnx so much!!

OpenStudy (mathstudent55):

wlcm

OpenStudy (destinymasha):

i gave u a medal right?

OpenStudy (mathstudent55):

yes, thanks